The following arbitrary measurements are made and the errors sited are the maximum errors A=15.21 +/- 0.01, B=10.82 +/-0.05, C=11.00 +/- 0.03. If D = A+B+C; (a) calculate the maximum error in D (b) If the errors sited are standard errors, calculate the standard error in D
1
Expert's answer
2021-08-16T08:37:04-0400

(a) To calculate the maximum error in D, we need to add all errors:


"\Delta_\text{max}D=0.01+0.05+0.3=0.09."

(b) Find the mean:


"\mu=\frac13(0.01+0.05+0.03)=0.03."

Find the standard error:


"s.e.=\\=\sqrt{(0.01-0.03)^2+(0.05-0.03)^2+(0.03-0.03)^2}=\\\space
=0.02."